# A Survey on Quantum Machine Learning: Basics, Current Trends, Challenges, Opportunities, and the Road Ahead

### v3

@@ 7.2 Quantum Sensing
GPS, radar, lidar, and other electromagnetic technologies use quantum physics to provide increasingly common tools on city streets. In aircraft and even on basic cell phones, quantum sensing is in the process of shifting from being a highly-prized capability few can afford to being in daily use everywhere. Once it achieves widespread adoption, quantum sensing is expected to improve capabilities for the following domains dramatically:
@@ 7.4 Financial Modeling
By enabling more accurate and complex financial data modeling, improving risk management, and optimizing investment portfolios, QML could lead to better and informed decision-making with higher returns for investors. It can also enable financial services organizations to re-engineer operational processes (Orús et al., 2019). QML’s specific use cases for financial services can be classified into three main categories: targeted predictions, trading optimizations, and risk profiling. Quantum technology computes more efficiently than classical computing when generating probability distributions, mapping data, testing samples, and iterating. These properties align well with the financial modeling since they can help enhance investment gains, reduce capital requirements, open new investment opportunities, and improve the identification and management of risk and compliance. The simplest instance of practical implementation of QML for finance can include a task of portfolio optimization using numerous quantum optimization algorithms, fundamentally derived from the parameterized quantum circuits. If not at the current stage, ultimately such tasks will benefit in terms of compute speed and processing information capacity for solutions using QML techniques.
@@ 7.5 Logistics Optimization
Supply chain and logistics professionals have been overloaded over the past several years. An increasing amount of uncertainties, extreme weather conditions, and pandemic-fueled supply and demand changes have exponentially increased logistics complexity. While businesses need supply chain optimizations to account for the complexity of the entire ecosystem, existing optimizations no longer provide the cure to supply chain and logistics challenges. In this regard, QML could provide solutions for persistent supply-chain and logistics problems, such as air traffic routing, trade shipment routing, telecom networks, quantum internet, and healthcare. Organizations in the transportation and logistics area could be among those to gain the most advantage from quantum computing’s groundbreaking capabilities. An organization’s path from the initial investigation of QML’s impact to actual readiness for implementation can span a few years.
@@ 7.6 Environmental, Social, and Governance (ESG)
With the current environmental situation, we see the global need to establish and ensure sustainability. Despite that, technological developments are no further uncorrelated to the environment. Hence, while the quantum technology itself should be designed to be environmentally sustainable, it can be a catalyst in enabling sustainable energy solutions at a large infrastructure scale. Quantum technology, apart from reducing the efficiency and execution time, can also introduce and develop methods that complement Environmental, Social, and Governance (ESG) goals. It can drastically reduce energy wastage and fuel costs with its potential in route optimizations and sensing technologies. Moreover, it provides solutions that empower people to leave aside tedious and repetitive tasks and has a huge social impact by fulfilling problems that drastically improve the quality of life. Therefore, there is great potential to evolve the quantum field from the ESG sustainability perspective.
@@ B.1 Hilbert Space
B.2 State Space and State Vector
@@ B.3 Schrodinger’s Equation
B.4 Hermitian Matrix/Operator
@@ B.5 Observable
B.6 Heisenberg’s Uncertainty Principle
This principle states that we cannot know both the position and speed of a particle, such as a photon or electron, with perfect accuracy. The more we accurately know the position of the particle, the less we know about its speed, and vice versa.

### v4

@@ 7 Applications
However, demonstrating a real-world quantum advantage in these application domains is contingent on further progress in quantum hardware and on the development of application-specific quantum architectures. In practice, quantum solutions will need more qubits with lower noise and longer coherence times to outperform classical methods on meaningful tasks. Therefore, it is essential to conduct a co-design approach in which quantum algorithms and hardware are engineered together for specific use cases. Such application-specific quantum computing architectures, tuned to particular problem requirements, are key for unlocking practical QML advantages. In other words, with the help of future hardware improvements and specialized quantum algorithm–hardware optimization, QML will make significant strides in real-world applications.
@@ 7.2 Quantum Sensing
GPS, radar, lidar, and other electromagnetic technologies use quantum physics to provide increasingly common tools on city streets. In aircraft and even on basic cell phones, quantum sensing is in the process of shifting from being a highly-prized capability few can afford to being in daily use everywhere. Once it achieves widespread adoption, quantum sensing is expected to significantly improve capabilities for the following domains:
@@ 7.4 Financial Modeling
Quantum-enhanced portfolio optimization and risk analysis have been demonstrated on small benchmark problems, showing that variational quantum algorithms can achieve comparable solutions to classical solvers with fewer iterations (Orús et al., 2019; Egger et al., 2020). For instance, the VQE and QAOA have been applied to relatively small portfolio instances, yielding near-optimal allocations in simulation (Havlíček et al., 2019; Zaman et al., 2024b). Likewise, quantum-inspired Monte Carlo methods leverage amplitude estimation to compute value at risk and option pricing with quadratic speedup in the fault-tolerant regime (Montanaro, 2015; Woerner & Egger, 2019). Moreover, hybrid quantum-classical ML systems have been proven successful for financial fraud detection prevention (Grossi et al., 2022; Innan et al., 2025).
@@ 7.5 Logistics Optimization
Supply chain and logistics professionals have been overloaded over the past several years. An increasing amount of uncertainties, extreme weather conditions, and pandemic-fueled supply and demand changes have exponentially increased logistics complexity. Combinatorial optimization in routing and scheduling can be framed as QUBO problems and solved via quantum annealing or QAOA. Early studies mapped small Vehicle Routing Problems (VRPs) to D-Wave hardware, achieving feasible tours but with limited scaling due to embedding overhead (Venturelli et al., 2015; Neukart et al., 2017). Variational algorithms on gate-model devices have tackled small Traveling Salesman Problem (TSP) instances, demonstrating proof-of-concept quantum speedups (Cerezo et al., 2021a). In practice, embedding real-world logistics networks (hundreds of nodes) remains out of reach on NISQ devices. Hence, hybrid workflows partition the problem into quantum-solvable subgraphs and classical refinement loops (Chicano et al., 2025).
@@ 7.6 Environmental, Social, and Governance (ESG)
Quantum sensing for environmental monitoring exploits enhanced sensitivity of entangled states. Proof-of-principle experiments using spin-squeezed Bose–Einstein condensates have measured magnetic fields with precision beyond the standard quantum limit (Pezze et al., 2018; Hosten et al., 2016). Such sensors could track pollutant concentrations or micro-climate variables with greater accuracy, benefiting ESG reporting. QML can support ESG goals by optimizing renewable-microgrid scheduling and emissions reduction. For instance, a hybrid quantum-classical QAOA approach to microgrid economic dispatch reported lower operational cost and resilience under fluctuating renewable supply compared to classical methods (Koretsky et al., 2021). In environmental monitoring, quantum-enhanced sensing algorithms using entangled probes have demonstrated improved precision in detecting trace gases (Degen et al., 2017). However, broad ESG impact hinges on scaling these methods beyond prototypical demos toward fault-tolerant platforms and integrating them into existing infrastructure.
@@ 8 Conclusion and Road Ahead
Finally, it is important to explicitly acknowledge broader ethical and societal impacts associated with QC and QML. The emergence of large-scale fault-tolerant quantum computers poses substantial threats to current cryptographic standards (e.g., RSA encryption), potentially undermining cybersecurity infrastructures and necessitating timely development and adoption of quantum-safe cryptographic methods. Moreover, access to quantum computing resources and expertise remains limited and costly, potentially exacerbating existing inequalities between institutions and nations with differing economic capabilities. The significant energy demands of current quantum systems, especially those reliant on cryogenic cooling, further highlight environmental sustainability concerns that must be addressed as this technology scales. Additionally, QML inherits ethical considerations common to classical AI/ML, such as data biases, fairness issues, transparency, interpretability challenges, and risks of misuse in surveillance or autonomous weapon systems. Addressing these ethical and societal challenges proactively will be essential for responsible and equitable advancement in quantum technologies.
@@ B.1 Hilbert Space
B.2 CHSH Inequality
The Clauser–Horne–Shimony–Holt (CHSH) inequality is a key result in quantum physics that distinguishes the predictions of quantum mechanics from classical physics, particularly those based on local hidden variable theories. In a CHSH experiment, two observers (Alice and Bob) make measurements on entangled particles using randomly chosen settings (Bell, 1964). Classical theories constrained by locality and realism predict a specific bound on the correlation between their outcomes, encapsulated by the CHSH inequality(Clauser et al., 1969).
Quantum mechanics, however, predicts stronger correlations that violate this classical bound when entangled particles are involved (Cirel’son, 1980). Repeated experiments have confirmed this violation, supporting quantum mechanics and challenging classical assumptions (Aspect et al., 1982; Hensen et al., 2015; Shalm et al., 2015; Giustina et al., 2015). These findings not only deepen our understanding of quantum entanglement but also form the basis for future quantum computing technologies.
B.3 Wave-Particle Duality
B.4 Max-Born Rule
B.5 State Space and State Vector
@@ B.6 Schrodinger’s Equation
B.7 Hermitian Matrix/Operator
@@ B.8 Observable
B.9 Heisenberg’s Uncertainty Principle
Quantum noise arises from imperfections in quantum operations and the measurement process, rather than from inherent randomness in quantum mechanics itself. Although quantum dynamics are fundamentally deterministic, practical quantum systems are subject to environmental interactions and operational inaccuracies, causing deviations from ideal quantum states (Nielsen & Chuang, 2010b). While the Heisenberg uncertainty principle limits the simultaneous precision of certain observables, it does not directly produce noise. In contrast, quantum noise is primarily due to errors introduced during state preparation, gate operations, and measurements, collectively impacting system fidelity and performance (Preskill, 2018).
$\Delta x\cdot\Delta p\geq\frac{\hbar}{2}$ (4)